Your firm has an average receipt size of $125. A bank has approached you concerning a lockbox service that will decrease your total collection time by two days. You typically receive 6,400 checks per day. The daily interest rate is 0.016 percent. The bank charges a lockbox fee of $175 per day.

What is the NPV of accepting the lockbox agreement?

What would the net annual savings be if the service were adopted? (Use 365 days a year. Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16))
Answer #1

NPV = 6400*125*2 -175/.00016 = 506,250

---

Annual savings = 6400*125*2*(1+0.00016)^365 - 6400*125*2 = 96214.42

Annual cost = 175*FVIFA(0.00016,365) = 175*((1+0.00016)^365-1)/0.00016 = 65,771.58

Annual Savings = 96214.42 - 65,771.58 = 30442.84